Phil had 50 nickels and dimes worth $4.15. How many dimes does he have.
Answer by Maths68(1474)

Total worth = $4.15 = 4.15*100 = 415 cents
Total Coins = 50
Total Dimes = x
Total Nickel = 50-x
One Dime = d = 10 cents
One Nickel = n = 5 cents
Dime*x+Nickel*(50-x)=Total worth
10*x+5(50-x)=415
10x+250-5x=415
5x=415-250
5x=165
5x/5=165/5
x=33


Total Dimes = x = 33
Total Nickel = 50-x = 50-33 = 17


Check
=========
10*33+5*17=415
330+85=415
415=415